OpenStudy (anonymous):

There were 150 people at the Junior-Senior dance Sasha and her friends attended. Junior tickets were $2.00 each and Senior tickets were $3.50 each. The total receipts for the dance were $405. How many Juniors bought tickets?

OpenStudy (mathstudent55):

Let x = number of junior Let y = number of seniors Then x + y = 150 Amount of money raised by selling x tickets to juniors: $2.00(x) = 2x Amount of money raised by selling y tickets to seniors:: $3.50(y) = 3.5y Total amount of money raised: 2x + 3.5y = 405 Solve the two equations simultaneously: x + y = 150 2x + 3.5y = 405